OpenStudy (anonymous):

sin(2Θ-π)= sinΘ SOLVE ON THE INTERVAL 0≤Θ<2π

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

i would start with \[\theta=20-\pi\]

OpenStudy (anonymous):

of course \(20-\pi\) is not in that interval so keep subtracting \(2\pi\) until it is

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

can you post a screen shot of the actual question?

OpenStudy (anonymous):

well would you 1st plug in double angle identity on left side

OpenStudy (anonymous):

its exactly what ive posted

OpenStudy (anonymous):

no, both \(20\) and \(\pi\) are numbers and so there is no "double angle" identity to use

OpenStudy (anonymous):

and sorry i meant sum and difference identity not double angle

OpenStudy (anonymous):

its not 20, its 2theta

OpenStudy (anonymous):

the only problem is that \(20-\pi\) in the interval \(0<\theta<2\pi\) it is way bigger

OpenStudy (anonymous):

ooooh!

OpenStudy (anonymous):

\[\sin(2\theta-\pi)=\sin(\theta)\] like that?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

ok now we can get somewhere

OpenStudy (anonymous):

ok so whatd be 1st step

OpenStudy (anonymous):

i would write \[\sin(2\theta-\pi)=-\sin(2\theta)\] first

OpenStudy (anonymous):

then use the double angle formula

OpenStudy (anonymous):

why negative

OpenStudy (anonymous):

same reason that \[\sin(x-\pi)=-\sin(x)\] if you look on the unit circle you will see why

OpenStudy (anonymous):

so 2sinΘcosΘ-1= -sin2Θ ?

OpenStudy (anonymous):

then the double angle formula for \(\sin(2\theta)\) gives \[-2\sin(\theta)\cos(\theta)=\sin(\theta)\]

OpenStudy (anonymous):

are u saying it could be either? or was i wrong?

OpenStudy (anonymous):

i think i might have lost you somewhere in there want to go slow?

OpenStudy (anonymous):

cuz this is what i got. 2sinΘcosΘ-1= -sin2Θ ?

OpenStudy (anonymous):

on the right hand side of the equal sign you have \(\sin(\theta)\) right?

OpenStudy (anonymous):

sin2Θ=2sinΘcosΘ thats the formula

OpenStudy (anonymous):

& yea i do

OpenStudy (anonymous):

so this is the original question \[\sin(2\theta-\pi)=\sin(\theta)\]

OpenStudy (anonymous):

but u said to make it -sin2Θ

OpenStudy (anonymous):

okay. then?

OpenStudy (anonymous):

lets leave the right hand side alone

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and work only with \[\sin(2\theta-\pi)\]

OpenStudy (anonymous):

so 1st step?

OpenStudy (anonymous):

rewrite \(\sin(2\theta-\pi)\) as \(-\sin(2\theta)\)

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

ok so π is gone?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so that equals right hand side. so far i get it

OpenStudy (anonymous):

because it is always the case that \[\sin(\xi-\pi)=-\sin(\xi)\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok now use the "double angle" formula for \(-\sin(2\theta)\)

OpenStudy (anonymous):

so itd be same formula only it would hav negative in front?

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

that is \[\sin(2\theta)=2\sin(\theta)\cos(\theta)\] so \[-\sin(2\theta)=-2\sin(\theta)\cos(\theta)\]

OpenStudy (anonymous):

-2sinΘcosΘ=cosΘ

OpenStudy (anonymous):

yeah, but you messed with the right hand side, it is still \(\sin(\theta)\)

OpenStudy (anonymous):

sorry i meant that

OpenStudy (anonymous):

ok good now set it equal to zero, by adding \(\sin(\theta)\cos(\theta)\) to both sides, giving \[2\sin(\theta)\cos(\theta)+\sin(\theta)=0\]

OpenStudy (anonymous):

ok so bring everything to one side. then what?

OpenStudy (anonymous):

factor out the common factor of \(\sin(\theta)\)

OpenStudy (anonymous):

when you factor what does it look like

OpenStudy (anonymous):

\[\sin(\theta)\left(2\cos(\theta)+1\right)=0\]

OpenStudy (anonymous):

and then solve for each part? set each equal to 0 and solve?

OpenStudy (anonymous):

set each factor equal to zero, solve for \(\theta\) that is, solve \[\sin(\theta)=0\]and solve \[2\cos(\theta)+1=0\]exactly, what you said

OpenStudy (anonymous):

so answers would be π, 2π/3 & 4π/3

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

is \(0\) in your interval?

OpenStudy (anonymous):

0 as well

OpenStudy (anonymous):

if the interval is \(0\leq \theta\leq 2\pi\) i.e. the closed interval, you have for sine 3 solutions \[0,\pi,2\pi\]

OpenStudy (anonymous):

no not 2π

OpenStudy (anonymous):

for cosine you have the right two solutions

OpenStudy (anonymous):

cuz it has to be less than 2π

OpenStudy (anonymous):

ok yw